OpenStudy (anonymous):

(4x2 - 9) divided by (2x - 3)

OpenStudy (jdoe0001):

would you say that \(\large 4x^2 = 2^2x^2 = (2x)^2 \ ?\)

OpenStudy (anonymous):

No, i would say.. Find the quotient and enter your answer in descending order below. Enter exponents using the caret ( ^ ). (4x^2 - 9)/(2x - 3)

OpenStudy (anonymous):

but if you wanna say it like that you can, lol. idk how to do it. :(

OpenStudy (jdoe0001):

well, would you rather have long division?

OpenStudy (anonymous):

yeah!

OpenStudy (jdoe0001):

ok, gimme a few

OpenStudy (anonymous):

take ya time. :)

OpenStudy (jdoe0001):